OpenStudy (anonymous):

Find the equation of a line parallel to the given line: y = -3/4x -1. y = -3/4x + 2 y = 3/4x – 1 y = 4/3x – 1 y = ¾ x + 1

OpenStudy (anonymous):

line parallel to a line parallel, will have the same slope. y=mx+b is the equation for a line where m is the slope. In the given problem y = -3/4x -1, slope of this line, m=-3/4 so answer will be y = -3/4x + 2